How many moles of oxygen (O2) are needed to react with 56.8 grams of ammonia?

4 NH3 + 5 O2 → 4 NO + 6 H2O 


1
Expert's answer
2021-05-06T07:37:56-0400

M(NH3) = 17.03 g/mol

n(NH3) "= \frac{m}{M} = \frac{56.8}{17.03} = 3.335 \;mol"

According to the reaction:

n(O2) = "\frac{5}{4}" n(NH3) = "\frac{5}{4} \times 3.335 = 4.17 \;mol"

Answer: 4.17 mol